# 360VFI: A Dataset and Benchmark for Omnidirectional Video Frame Interpolation

### v2

@@ 4.2.1 DistortionGuard: A Distortion-Aware Pyramid Feature Extractor
Previous methods tend to treat $C_{d}$ as an additional input of $X$ (Nishiyama et al., 2021), or re-weighting parameters by $C_{d}$ (Khasanova & Frossard, 2019). However, continuous and amorphous distortions cannot be adequately fitted by scattering and structured convolution operations. Therefore, we intend to design a novel block to continuously learn distorted patterns. In video frame interpolation tasks, the deformable mechanism is proposed to align features between adjacent frames (Tian et al., 2020; Wang et al., 2020b). Unlike standard DCN (Dai et al., 2017), which calculates offsets from the input feature map, offsets are calculated from bi-directional optical flow in VSR pipelines.
@@ 4.2.2 OmniFTB: Omnidirectional Distortion Feature Transform Block
Figure 7: OmniFTB (A Distortion-aware Frame Generator): (a) shows the details of OmniFTB that is composed of Distortion-aware Feature Transform (DFT) layers and other parts; (b) shows the details of a DFT layer using affine transformation. It transforms the feature $\bm{F}^{*}$ from the last convolution layer with the parameter from the ERP distortion map into the feature with distortion $\bm{F}^{***}$. Hence DFT layers recover ERP distortion from less distorted features.
Unlike DistortionGuard, which uses priors in the feature extraction stage, the affine transformation is an efficient way to perform feature-wise transformations (Dumoulin et al., 2018). Consequently, in the intermediate feature reconstruction of our method, we propose another block considering ERP priors.
For the feature output from the extraction stage, the distortion is modulated successfully. So we devise a distortion-aware feature transform block (OmniFTB) to apply an affine transformation with a pair of distortion-based parameters. In OmniFTB, a mapping function $M_{\theta^{\prime}}$ based on ERP distortion condition map $\bm{C_{d}}$ is learned at first,
@@
where $\bm{F}^{***}$ denotes the intermediate feature maps, whose dimension is the same as $\bm{\alpha}$ and $\bm{\beta}$, and $\odot$ is referred to element-wise multiplication, i.e., Hadamard product. Since the spatial dimensions are preserved, the SFT layer not only performs feature-wise manipulation but also spatial-wise transformation. Figure 7. (b)shows an example of implementing DFT block embedded in the intermediate frame generating decoder. The mapping function $M_{\theta^{\prime}}$ can be arbitrary functions. In this study, we use a neural network for $M_{\theta^{\prime}}$ so that it can be optimized end-to-end with the Omni-VFI branch. OmniFTB complements DistortionGuard by focusing on intermediate feature reconstruction. By applying an affine transformation with distortion-based parameters, OmniFTB effectively leverages ERP distortion priors to transform feature maps for accurate intermediate frame generation.
@@ 5.1 Implementation Details
We have implemented the proposed algorithm in PyTorch and utilized the 360VFI dataset to train 360VFI Net from scratch. Our model is optimized by AdamW (Loshchilov & Hutter, 2019) algorithm for 300 epochs with a total batch size of 8 on Two NVIDIA Tesla V100 GPUs. Initially, the learning rate is set to $1\times 10^{-4}$ and gradually decays to $1\times 10^{-5}$ following a cosine attenuation schedule. Throughout the training process, we refrained from applying augmentation techniques such as rotating and random cropping to triplet samples. This decision was made due to the rigid nature of latitude-dependent distortion.
@@ 5.2 Evaluation and Comparison
We conducted comparative experiments to evaluate the performance of 360VFI Net against existing methods. We utilized the first omnidirectional video dataset 360VFI we proposed and compared our method with several common VFI approaches. We use the WSS-L1 Loss (Baniya et al., 2023) as the loss function to train 360VFI Net. Smooth L1 loss acts as both L1 and L2 losses conditioned to a hyper-parameter $\beta$. It combines the advantages of L1-loss (steady gradients for large values) and L2-loss (less oscillation during update when values are small). Thus, it is less sensitive to outliers and prevents exploding gradients in some cases. However, the loss functions used for conventional images and videos do not take into account the unique nature of ERP frames. Distortion across latitude present in ERP frames can cause learning-based models to be easily influenced by high errors in prediction across polar regions. Therefore, following in the footsteps of WS-PSNR, a Weighted Spherically Smooth-L1 (WSS-L1) Loss function was proposed in Baniya et al. (2023) for taking distortion into consideration along with Smooth-L1 loss,
@@
where $\psi$ is the ERP distortion map and $\beta$ is a hyper-parameter.
We compared the proposed 360VFI Net module with the following methods: IFRNet, DQBC, EMA-VFI, EBME and UPR-Net. We employed the following evaluation metrics to assess the performance of different interpolation methods: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), Weighted PSNR (WS-PSNR) (Sun et al., 2017b) and Weighted Structural Similarity Index (WS-SSIM) (Zhou et al., 2018). WS-PSNR extends PSNR by considering the importance of different regions in ODV, calculating PSNR values with weighting factors to emphasize regions of interest, such as central regions with high visual significance. Similarly, WS-SSIM extends SSIM by incorporating weighted factors to account for the perceptual importance of different regions in ODV. It provides a more accurate assessment of perceptual quality considering distortion. With the comprehensive evaluation metrics, we aim to thoroughly evaluate the performance of Omni-VFI methods on four settings of different motion extents, considering both fidelity to ground truth and perceptual quality across the entire omnidirectional view.
@@ 5.3 Ablation Study
To validate the effectiveness of the two key modules, DistortionGuard and OmniFTB, within 360VFI Net, we conducted ablation experiments. In these experiments, we systematically removed each module and evaluated its impact on the final performance. Specifically, in the DistortionGuard module ablation experiment, we constructed a variant of 360VFI Net without the DistortionGuard module and compared it to the complete 360VFI Net. Similarly, in the OmniFTB module ablation experiment, we designed a variant of 360VFI Net without the OmniFTB module and contrasted it with the version containing the complete architecture. We maintained the same dataset and training settings and compared the performance of the two versions. Table 4 shows that both ablated versions exhibited a noticeable drop in performance compared to the complete 360VFI Net, highlighting the importance of both DistortionGuard and OmniFTB in achieving optimal results.
The results of the ablation experiments demonstrate the significant impact of the DistortionGuard and OmniFTB modules on the performance of 360VFI Net. Their presence enhances the model’s robustness and accuracy, providing crucial support for Omnidirectional Video Frame Interpolation tasks. These findings further substantiate the effectiveness and reliability of 360VFI Net in handling omnidirectional videos, laying a solid foundation for future research and applications.
@@ 6 Conclusion
In this paper, we have devised a dataset and benchmark termed 360VFI for Omnidirectional Video Frame Interpolation. We provide the first dataset for Omnidirectional Video Frame Interpolation, which is collected and cleaned from multiple sources and tailored into a triplet format. 360VFI dataset has four evaluation settings, serving as a benchmark for various extents of motion in Omnidirectional Videos and facilitating future research in this field. Furthermore, the proposed 360VFI Net uniquely incorporates ERP distortion priors in both the feature extraction stage and the frame generation stage, employing different customized methods for each stage. It first extracts less distorted features of two input frames according to the distortion map and then gradually generates the target ERP frame using parameterized affine transformation to recover the distortion. We anticipate that our contributions will inspire further advancements in omnidirectional video processing techniques, ultimately enhancing the immersive viewing experience.

### v3

@@ 4.2.1 DistortionGuard: A Distortion-Aware Pyramid Feature Extractor
Previous methods tend to treat $C_{d}$ as an additional input of $X$ (Nishiyama et al., 2021), or re-weighting parameters by $C_{d}$ (Khasanova & Frossard, 2019). However, continuous and amorphous distortions cannot be adequately fitted by scattering and structured convolution operations. Therefore, we intend to design a novel block to learn distorted patterns continuously. In video frame interpolation tasks, the deformable mechanism is proposed to align features between adjacent frames (Tian et al., 2020; Wang et al., 2020b). Unlike standard DCN (Dai et al., 2017), which calculates offsets from the input feature map, offsets are calculated from bi-directional optical flow in VSR pipelines.
@@ 4.2.2 OmniFTB: Omnidirectional Distortion Feature Transform Block
Figure 7: OmniFTB is a Distortion-aware Frame Generator: (a) shows the details of OmniFTB that are composed of Distortion-aware Feature Transform (DFT) layers and other parts; (b) shows the details of a DFT layer using affine transformation. It transforms the feature $\bm{F}^{*}$ from the last convolution layer with the parameter from the ERP distortion map into the feature with distortion $\bm{F}^{***}$. Hence, DFT layers recover ERP distortion from less distorted features.
Unlike DistortionGuard, which uses priors in the feature extraction stage, the affine transformation efficiently performs feature-wise transformations (Dumoulin et al., 2018). Consequently, in the intermediate feature reconstruction of our method, we propose another block considering ERP priors.
For the feature output from the extraction stage, the distortion is modulated successfully. So, we devise a distortion-aware feature transform block (OmniFTB) to apply an affine transformation with a pair of distortion-based parameters. In OmniFTB, a mapping function $M_{\theta^{\prime}}$ based on ERP distortion condition map $\bm{C_{d}}$ is learned at first,
@@
where $\bm{F}^{***}$ denotes the intermediate feature maps, whose dimension is the same as $\bm{\alpha}$ and $\bm{\beta}$, and $\odot$ is referred to element-wise multiplication, i.e., Hadamard product. Since the spatial dimensions are preserved, the SFT layer not only performs feature-wise manipulation but also spatial-wise transformation. Figure 7. (b)shows an example of implementing DFT block embedded in the intermediate frame generating decoder. The mapping function $M_{\theta^{\prime}}$ can be arbitrary functions. In this study, we use a neural network for $M_{\theta^{\prime}}$ to optimize it end-to-end with the Omni-VFI branch. OmniFTB complements DistortionGuard by focusing on intermediate feature reconstruction. By applying an affine transformation with distortion-based parameters, OmniFTB effectively leverages ERP distortion priors to transform feature maps for accurate intermediate frame generation.
@@ 5.1 Implementation Details
We implement the proposed algorithm in PyTorch and utilize the 360VFI dataset to train the 360VFI network from scratch. Our model is optimized by AdamW (Loshchilov & Hutter, 2019) algorithm for 300 epochs with a total batch size of 8 on Two NVIDIA Tesla V100 GPUs. Initially, the learning rate is set to $1\times 10^{-4}$ and gradually decays to $1\times 10^{-5}$ following a cosine attenuation schedule. Throughout the training process, we refrained from applying augmentation techniques such as rotating and random cropping to triplet samples. This decision was made due to the rigid nature of latitude-dependent distortion.
@@ 5.2 Evaluation and Comparison
We conducted comparative experiments to evaluate the performance of 360VFI Net against existing methods. We utilized the first omnidirectional video dataset 360VFI we proposed and compared our method with several common VFI approaches. We adopt the WSS-L1 Loss (Baniya et al., 2023) as the loss function for training 360VFI Net. The Smooth L1 loss combines the strengths of both L1 and L2 losses, governed by a hyper-parameter $\beta$. It offers the benefits of L1 loss (steady gradients for large errors) and L2 loss (less oscillation for small errors), making it more robust to outliers and preventing exploding gradients in certain scenarios. However, traditional loss functions do not account for the unique characteristics of ERP frames, particularly the distortion that occurs across latitudes. This distortion can lead to significant prediction errors, especially in polar regions, when using learning-based models.
To address this issue, the Weighted Spherically Smooth-L1 (WSS-L1) Loss (Baniya et al., 2023) was introduced, building on the idea of WS-PSNR. The WSS-L1 Loss adjusts the Smooth L1 loss to account for ERP distortion, as shown in the following equation:
@@
Here, $\psi$ represents the ERP distortion map, and $\beta$ is the hyper-parameter that controls the transition between L1 and L2 behavior.
We compared the proposed 360VFI Net module with the following methods: IFRNet (Kong et al., 2022), DQB (Zhou et al., 2023), EMA-VFI (Zhang et al., 2023), EBME (Jin et al., 2023b) and UPR-Net (Jin et al., 2023a). We employed the following evaluation metrics to assess the performance of different interpolation methods: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), Weighted PSNR (WS-PSNR) (Sun et al., 2017b) and Weighted Structural Similarity Index (WS-SSIM) (Zhou et al., 2018). WS-PSNR extends PSNR by considering the importance of different regions in ODV and calculating PSNR values with weighting factors to emphasize regions of interest, such as central regions with high visual significance. Similarly, WS-SSIM extends SSIM by incorporating weighted factors to account for the perceptual importance of different regions in ODV. It provides a more accurate assessment of perceptual quality considering distortion. With the comprehensive evaluation metrics, we aim to thoroughly evaluate the performance of Omni-VFI methods on four settings of different motion extents, considering both fidelity to ground truth and perceptual quality across the entire omnidirectional view.
@@ 5.3 Ablation Study
To assess the effectiveness of the two key components, DistortionGuard and OmniFTB, within the 360VFI Net architecture, we performed a series of ablation experiments. In these experiments, we systematically removed each module and evaluated its impact on overall performance. Specifically, for the DistortionGuard ablation, we created a variant of 360VFI Net without the DistortionGuard module and compared its performance against the total 360VFI Net. Similarly, for the OmniFTB ablation, we constructed a variant of the model without the OmniFTB module and contrasted it with the complete architecture.
Both experiments were conducted under identical datasets and training conditions. As shown in Table 4, the removal of either module resulted in a significant performance degradation compared to the full model, underscoring the critical role that both DistortionGuard and OmniFTB play in achieving optimal results.
The ablation results clearly demonstrate the substantial contribution of the DistortionGuard and OmniFTB modules to the overall performance of the 360VFI network. Their inclusion enhances the model’s robustness and accuracy, playing an essential role in omnidirectional video frame interpolation tasks. These findings further validate the effectiveness and reliability of the 360VFI Net architecture, providing a solid foundation for future research and development in this domain.
@@ 6 Conclusion
360VFI is the first dataset and benchmark for omnidirectional video frame interpolation. Our dataset has four evaluation settings, serving as a benchmark for various extents of motion in omnidirectional videos and facilitating future research in this field. Furthermore, the proposed 360VFI network uniquely incorporates ERP distortion priors in both the feature extraction stage and the frame generation stage, employing different customized methods for each stage. It first extracts less distorted features of two input frames according to the distortion map and then gradually generates the target ERP frame using parameterized affine transformation to recover the distortion. We anticipate our contributions will inspire further advancements in omnidirectional video processing techniques, ultimately enhancing the immersive viewing experience.